### v1

# Entity or Relation Embeddings? An Analysis of Encoding Strategies for Relation Extraction

Abstract
Relation extraction is essentially a text classification problem, which can be tackled by fine-tuning a pre-trained language model (LM). However, a key challenge arises from the fact that relation extraction cannot straightforwardly be reduced to sequence or token classification. Existing approaches therefore solve the problem in an indirect way: they fine-tune an LM to learn embeddings of the head and tail entities, and then predict the relationship from these entity embeddings. Our hypothesis in this paper is that relation extraction models can be improved by capturing relationships in a more direct way. In particular, we experiment with appending a prompt with a [MASK] token, whose contextualised representation is treated as a relation embedding. While, on its own, this strategy significantly underperforms the aforementioned approach, we find that the resulting relation embeddings are highly complementary to what is captured by embeddings of the head and tail entity. By jointly considering both types of representations, we end up with a simple model that outperforms the state-of-the-art across several relation extraction benchmarks.
Entity or Relation Embeddings?
An Analysis of Encoding Strategies for Relation Extraction
Frank Mtumbuka and Steven Schockaert Cardiff University, UK {MtumbukaF,SchockaertS1}@cardiff.ac.uk
1 Introduction
Relation extraction consists in identifying the relationships between entities that are expressed in text. It is a fundamental Natural Language Processing task (NLP), which enables the learning of symbolic representations such as knowledge graphs. While Large Language Models (LLM) are highly effective in interpreting natural language inputs, the state-of-the-art in relation extraction is still based on fine-tuned language models of the BERT family Devlin et al. (2019). Moreover, relation extraction is often applied to large document collections, such as corpora of news stories, scientific articles or social media, which means that relying on LLMs is often not feasible in practice, due to their high cost.
In high-level terms, the standard approach is to fine-tune a BERT-based language model to learn a relation embedding, i.e. a vector representation of the relationship that is expressed between two entities in a given sentence, and to train a classifier to predict a discrete relation label from this embedding. To obtain these relation embeddings, we cannot simply train a sentence embedding model, since a single sentence may express several relationships. Moreover, we cannot easily identify which tokens in the sentence express the relationship, which makes learning relation embeddings fundamentally different from learning embeddings of text spans. In their seminal work, Baldini Soares et al. (2019) proposed to encapsulate the head and tail entity with special tokens. Consider the following example, where we are interested in extracting the relationship between Paris and France:
The Olympics will take place in [E1] Paris [/E1], the capital of [E2] France [/E2].
The corresponding relation embedding is obtained by concatenating the final-layer embedding of the special tokens [E1] and [E2]. This strategy was found to outperform other alternatives by Baldini Soares et al. (2019), and has remained the most popular approach for learning relation embeddings (although Zhou and Chen (2022), among others, have used type-specific special tokens, which encode the semantic type of the entities).
The success of this strategy is somewhat surprising: [E1] and [E2] are intuitively designed to represent the head and tail entities, rather than their relationship. In practice, however, knowing the semantic types of the entities often allows us to “guess” the relationship between them, especially if the types are fine-grained. For instance, knowing that the head entity is a capital city and the tail entity is a country, we may reasonably assume that the relationship being expressed is capital of. Nonetheless, we can expect that approaches based on entity embeddings may struggle to distinguish between closely related relations, especially if these relations involve entities of the same type. A possible alternative is to add a prompt which includes the [MASK] token, e.g.:
The Olympics will take place in Paris, the capital of France. The relation between Paris and France is [MASK].
We can then simply fine-tune the LM such that the contextualised embedding of the [MASK] token corresponds to a relation embedding. This strategy is popular for zero-shot and few-shot relation extraction Genest et al. (2022), but is not normally used for the standard supervised relation extraction setting. Our experiments indeed confirm that this strategy, on its own, performs poorly. Our hypothesis is that this strategy is hampered by the fact that it cannot adequately model the semantic types of the entities, which makes relation prediction much harder. The limitations of this strategy are thus complementary to those of the more common entity embedding based strategy: in one case the model almost exclusively focuses on entity types while in the other case it largely ignores them.
We present an experimental analysis which supports this hypothesis. In particular, we find that by simply combining both strategies, we can achieve substantial performance improvements, allowing us to improve the state-of-the-art in several relation extraction benchmarks. We furthermore experiment with the use of a pre-trained entity embedding model Mtumbuka and Schockaert (2023), which, despite not being trained on the relation extraction task, perform as well as the approach from Baldini Soares et al. (2019), which clearly supports our main hypothesis. As a final contribution, we also propose the use of a self-supervised pre-training strategy, which brings further performance gains.
2 Related Work
Learning Relation Embeddings
The standard approach for relation extraction with LMs uses special tokens to indicate the head and tail entities (also known as the subject and object). Such approaches compute relation embeddings by concatenating the contextualised representations of these two entities. The relation label itself is then typically predicted by training a simple classifier that takes this relation embedding as input. Several variants of this approach have been considered. The matching-the-blanks model Baldini Soares et al. (2019) encapsulates the head entity using the special tokens [E1][/E1] and the tail entity using separate special tokens [E2][/E2]. LUKE Yamada et al. (2020) simply replaces the entities by the special tokens [HEAD] and [TAIL], omitting the actual entity spans from the input. Another possibility is to use special punctuation. For instance Wang et al. (2021) encapsulate the head and tail entity with the markers @…@ and #…#, thus avoiding the introduction of new tokens. More recently, promising results have been obtained by using typed markers, which encode the semantic type of the entity. Zhong and Chen (2021) simply use type-specific special tokens, e.g. encapsulating the head entity using tokens of the form [S:type][/S:type]. However, this requires the in introduction of a large number of special tokens. Zhou and Chen (2022) avoid this by using punctuation and verbalising the semantic type as part of input. For instance, the head entity is encapsulated as @ * type *@, with type a description of the entity type.
Another possibility is to append the input with a prompt containing the [MASK] token. This strategy is popular in the context of zero-shot and few-shot relation extraction Gong and Eldardiry (2021); Chen et al. (2022); Genest et al. (2022). Rather than training a classifier on top of a relation embedding, the main aim is then to compare the contextualised representation of the MASK token with verbalisers, i.e. tokens from the language model’s vocabulary that describe the relationship. However, to the best of our knowledge, this strategy has not yet been considered in the fully supervised setting.
Zhong and Chen (2021) already discussed the idea that representing entity types is not sufficient for relation classification. They highlight in particular that jointly training an entity typing and relation extraction system hurts performance, suggesting that both tasks need different kinds of latent representations. We take this idea further, based on the hypothesis that representing relations by concatenating the embeddings of the head and tail entity is inherently limited, even if the entity encoders are specifically trained for relation extraction.
Pre-training Relation Encoders
Several approaches have been proposed for pre-training or adapting language models to make them more suitable for the task of relation extraction. The matching-the-blanks model Baldini Soares et al. (2019) uses entity linking to find sentences that refer to the same entities, and then pre-trains a relation encoder based on the idea that sentences mentioning the same entity pairs are likely to express the same relationship. More recently, several authors have studied variants of this approach based on distant supervision. Two sentences are then assumed to have the same relation if their entity pairs are asserted to have the same relation in some knowledge base, a strategy that has a long tradition in relation extraction Mintz et al. (2009). For instance, Peng et al. (2020) implement this strategy using contrastive learning with the InfoNCE loss. While the aforementioned approaches are focused on fine-tuning a pre-trained LM, the idea of changing the LM model itself has also been explored. For instance, SpanBERT Joshi et al. (2020) changes the standard masked token prediction task with the aim of learning better span-level representations, while LUKE Yamada et al. (2020) uses entity linking and distinguished entity tokens to improve the representation of entities.
Relation Extraction with LLMs
Most work on relation extraction uses fine-tuned models of the BERT family Devlin et al. (2019). Wan et al. (2023) discuss some of the challenges involved in using LLMs for this task, including the difficulty in selecting suitable in-context demonstrations. For this reason, the focus is still on LMs that can be fine-tuned, although recent work has started to look at models that are (considerably) larger than the standard BERT-base and BERT-large models. For instance, Sainz et al. (2021) proposed a reformulation of relation extraction as a Natural Language Inference (NLI) problem and included experiments with the 1.5B parameter DeBERTaXXL model He et al. (2021). While this allowed them to improve the state-of-the-art at the time, it should be noted that the NLI based formulation is highly inefficient when a large number of relation labels need to be considered. It also relies on manually defined verbalisations of the relation labels. Wang et al. (2022) fine-tuned a 10B parameter model on a range of tasks that can be formulated as triple prediction, including relation extraction.
3 Relation Extraction
We consider the standard sentence-level relation extraction setting, where we are given a sentence in which two entities are highlighted, which we refer to as the head entity and tail entity. The goal is to predict which relationship holds between these two entities, given a pre-defined set of candidate relation labels. We focus on strategies that first learn a relation embedding, which describes the relationship between the two entities in a continuous space, and then use a classifier to predict the actual label based on that relation embedding. In Section 3.1 we first explain the pre-training strategies that we use for learning high-quality relation embeddings. Section 3.2 then describes how the pre-trained relation encoder is fine-tuned for the relation classification task. Finally, in Section 3.3 we explain how relation embeddings can be obtained by concatenating the contextualised embeddings of the head and tail entities or by using a prompt-based strategy with a [MASK] token, among others.
3.1 Pre-training the Relation Encoder
Pre-Training Objective
To pre-train relation encoders, we rely on the InfoNCE contrastive loss van den Oord et al. (2018), which has been found effective for learning relation embeddings Peng et al. (2020), and for representation learning in NLP more generally Gao et al. (2021); Liu et al. (2021); Li et al. (2023); Mtumbuka and Schockaert (2023). Specifically, let us assume that we have a set $S$ of sentences with designated head and tail entities. For each $s\in S$, we assume that we have access to a set of positive examples $P_{s}$, i.e. sentences which express the same relationship as the one expressed in $s$, and a set of negative examples $N_{s}$. Let us write $\phi(s)$ for the relation embedding obtained from sentence $s$ using some encoding strategy. For instance, $\phi(s)$ may be the concatenation of the contextualised representations of the head and tail entity, or it may be the representation of the [MASK] token when a relation prompt is used. Section 3.3 will describe the specific encoding strategies that we consider in our analysis. We train the encoder $\phi$ using the InfoNCE loss:
$-\sum_{s\in S}\sum_{p\in P_{s}}\log\frac{\exp\big{(}\cos(\phi(s),\phi(p))/\tau\big{)}}{\sum_{x}\exp\big{(}\cos(\phi(s),\phi(x))/\tau\big{)}}$ (1)
where the temperature $\tau>0$ is a hyperparameter, and the summation in the denominator ranges over $x\in N_{s}\cup\{p\}$. The loss captures the intuitive idea that two sentences that express the same relationship should be modelled using similar relation embeddings. As suggested by previous work Baldini Soares et al. (2019); Peng et al. (2020) we also include the masked language modelling objective during pre-trainig to prevent catastrophic forgetting. The overall loss is thus given by $\mathcal{L}_{\text{info}}+\mathcal{L}_{\text{MLM}}$, with $\mathcal{L}_{\text{info}}$ the loss in (1) and $\mathcal{L}_{\text{MLM}}$ the masked language modelling objective.
Self-Supervised Pre-Training
The effectiveness of the pre-training objective crucially depends on the quality and quantity of the available examples. In most cases, we have access to a set of labelled examples, obtained through manual annotation or distant supervision. The positive examples $P_{s}$ are then simply those examples that have the same label as $s$. As an alternative, we will also experiment with a form of self-supervised pre-training, using coreference chains as a supervision signal. Specifically, we adapt the EnCore strategy from Mtumbuka and Schockaert (2023) for pre-training entity encoders, by proposing a similar strategy for learning relation embeddings. The central idea is that two sentences are likely to express the same relationship if they refer to the same two entities. The matching-the-blank model also relies on this idea, but uses entity linking to identify such sentence pairs. Following Mtumbuka and Schockaert (2023) we select positive examples from the Gigaword corpus [^1] https://catalog.ldc.upenn.edu/LDC2003T05 and we only consider two entities to be co-referring if they are identified as such by two separate off-the-shelf coreference systems: the Explosion AI system Coreferee v1.3.1222https://github.com/explosion/coreferee and the AllenNLP coreference model [^3] https://demo.allennlp.org/coreference-resolution. This use of two coreference systems was found to reduce the number of false positives because of spurious coreference links, given that state-of-the-art coreference resolution systems are still far from perfect.
3.2 Relation Classification
Given a sentence $s$ with designated head and tail entities, we use the pre-trained relation encoder to obtain an embedding $\phi(s)$ that captures the relationship between these entities. Now consider the problem of classifying this relationship, using the labels from some set $\{l_{1},...,l_{m}\}$. Following standard practice Zhong and Chen (2021); Zhou and Chen (2022), we use a feedforward network with one hidden layer for this purpose:
$\mathbf{h}$ $=\mathsf{ReLU}(\mathbf{A_{1}}\phi(s)+\mathbf{b_{1}})$
$(p_{1},...,p_{m})$ $=\mathsf{softmax}(\mathbf{A_{2}}\mathbf{h}+\mathbf{b_{2}})$
where $p_{i}$ is interpreted as the probability that that $l_{i}$ is the correct label, $\mathbf{A_{1}}$ and $\mathbf{A_{2}}$ are matrices, and $\mathbf{b_{1}}$ and $\mathbf{b_{2}}$ are bias terms. The label classifier is trained using cross-entropy. We also fine-tune the pre-trained relation encoder during this step.
3.3 Encoding Strategies
We now discuss the strategies we consider for obtaining relation embeddings. Suppose we are interested in the relationship between the entities <h> and <t> in sentence $s$. We first create an annotated version of sentence $s$, where (i) <h> is encapsulated with the special tokens [E1][/E1] and <t> is encapsulated with the special tokens [E2][/E2], and we append the phrase “The relation between <h> and <t> is [MASK]”. For instance, for the example from the introduction we obtain the following annotated sentence:
The Olympics will take place in [E1] Paris [/E1], the capital of [E2] France [/E2]. The relation between Paris and France is [MASK].
For some of our strategies the [MASK] token will not be used, while another strategy only uses the [MASK] token. However, we use the same annotated sentence in all cases, as this allows for the most direct comparison. First, we consider the following basic approaches:
[H,T] We define $\phi(s)$ as the concatenation of the final-layer embeddings of the tokens [E1] and [E2], following Baldini Soares et al. (2019).
Mask We define $\phi(s)$ as the final-layer representation of the [MASK] token.
[H,T,Mask] We define $\phi(s)$ as the concatenation of the final-layer embeddings of the tokens [E1], [E2] and [MASK].
In each of the above cases, we pre-train the relation encoder as explained in Section 3.1 and then train a classifier as explained in Section 3.2.
Hybrid Strategy
We also consider a hybrid strategy, in which two different encoders are used during pre-training: $\phi_{1}$, which uses the [H,T] strategy, and $\phi_{2}$, which uses the Mask strategy. The pre-training loss is then of the form $\mathcal{L}_{\text{info}}^{1}+\mathcal{L}_{\text{info}}^{2}+\mathcal{L}_{\text{MLM}}$, where $\mathcal{L}_{\text{info}}^{1}$ and $\mathcal{L}_{\text{info}}^{2}$ both correspond to the InfoNCE loss (1), but where the former refers to the encoder $\phi_{1}$ and the latter refers to the encoder $\phi_{2}$. The encoders $\phi_{1}$ and $\phi_{2}$ share the same parameters, i.e. we fine-tune a single language model, but either use the concatenation of the embeddings of [E1] and [E2] as the relation embedding, or the embedding of the [MASK] token. After the pre-training step, when training the classifier, we simply concatenate the embeddings of the [E1], [E2] and [MASK] tokens. We will refer to this strategy as [H,T]+Mask. Note that the only difference with [H,T,Mask] lies in how the encoder is pre-trained. The motivation for the [H,T]+Mask strategy comes from the idea that the [MASK] token may be largely ignored when using the [H,T,Mask] strategy, since it may be easier to learn meaningful representations of the entities than to learn meaningful relation embeddings. The hybrid training strategy [H,T]+Mask avoids this issue, by forcing the [MASK] token to capture a meaningful relation embedding, before combining this representation with the contextualised entity embeddings.
Pre-trained Entity Embeddings
Our main hypothesis is that the [H,T] strategy focuses on learning the semantic types of the head and tail entities. To test this hypothesis, we consider a variant in which we use an entity embedding model instead of pre-training a relation encoder using the [H,T] strategy. In particular, we rely on EnCore Mtumbuka and Schockaert (2023) as the pre-trained entity embedding model. The EnCore embeddings essentially capture the semantic types of the entities (but without reference to any particular set of type labels). Crucially, no relational knowledge is used during the EnCore training process. We consider the following variants:
EnCore
No pre-training is used. We directly train a classifier on the concatenation of the EnCore embeddings of the head and tail entities.
EnCore+Mask
We pre-train an encoder using the Mask strategy. Then we train the classifier on the concatenation of the [MASK] token and the EnCore entity embeddings.
EnCore+[H,T]+Mask
We use the same hybrid pre-training as the [H,T]+Mask strategy. The classifier is then trained on the concatenation of the [H,T]+Mask representation and the EnCore entity embeddings.
Note that the EnCore model itself is not fine-tuned. This ensures that the embeddings provided by this model remain focused on entity types.
4 Experiments
In this section, we compare the effectiveness of the considered relation embedding strategies on a number of standard relation extraction benchmarks. Our main hypothesis is that the common [H,T] strategy essentially leads models to focus on the semantic types of the head and tail entity. We are thus interested in comparing the [H,T] and EnCore strategies. Furthermore, we hypothesise that the MASK strategy leads to embeddings capturing information that is complementary to that captured by the [H,T] embeddings. Accordingly, we are interested in comparing [H,T] with [H,T,Mask] and [H,T]+Mask. We first present the details of our experimental setup in Section 4.1. Section 4.2 subsequently summarises our main results. Finally, we present some additional analysis in Section 4.3.
4.1 Experimental Setup
Benchmarks
We consider a number of standard benchmarks for relation extraction. First, we use TACRED Zhang et al. (2017), a popular relation extraction benchmark. Two revisions of this dataset have been proposed, both of which are aimed at addressing some of the issues that arose in the original dataset due to the use of crowd-sourcing. First, TACREV Alt et al. (2020) was obtained by re-annotating the 5000 most challenging instances from the development and test sets. Specifically, to select these instances, the authors looked at how often models disagreed with each other and with the ground truth. Re-TACRED Stoica et al. (2021) was obtained by re-annotating the entire dataset. Following the tradition in recent work on relation extraction, we report results on all three variants in terms of F1 score. Next, we evaluate on a distantly supervised dataset that was introduced by Sorokin and Gurevych (2017) by aligning Wikipedia and Wikidata (Wiki-WD). Finally, we will also consider the distantly supervised dataset that was introduced by Riedel et al. (2010) by aligning articles from the New York Times with Freebase (NYT-FB). Following the tradition from previous work, we report the results on Wiki-WD in terms of F1 score and the results on NYT-FB in terms of precision at 10 (P@10) and 30 (P@30), averaged across all relation labels. Table 1 summarises the main characteristics of the considered datasets.
Dataset # Class Train Dev. Test
TACRED 42 68.1K 22.6K 15.5K
TACREV 42 68.1K 22.6K 15.5K
ReTACRED 40 58.4K 19.5K 13.4K
Wiki-WD 353 372.1K 123.8K 360.3K
NYT-FB 53 455.8K - 172.4K
Table 1: Overview of the considered benchmarks, showing the number of distinct relation classes, and the number of annotated mentions in the training, development and test sets.
Baselines
FOR TACRED and its variants, we consider a number of recent baselines. First, we include a comparison with SpanBERT Joshi et al. (2020) and LUKE Yamada et al. (2020). Next, we consider the Typed Marker strategy from Zhou and Chen (2022), and the Curriculum Learning variant from Park and Kim (2021). Note, however, that these last two methods are not directly comparable with our methods, as they rely on the gold entity type labels that are provided as part of the TACRED dataset. We do not use these labels for our models since such information is typically not available in practice. For Wiki-WD and NYT-FB, we compare with RECON Bastos et al. (2021) and KGPool Nadgeri et al. (2021). Note, however, that these state-of-the-art methods are again not directly comparable with our methods. They are focused on modelling and exploiting knowledge from the Wikidata and Freebase knowledge graphs. We do not take such information into account, as our focus is on comparing different encoding strategies for learning relation embeddings.
Training Details
Unless stated otherwise, the relation encoder is first pre-trained using the strategy from Section 3.1, before being fine-tuned, as explained in Section 3.2. In our default setting, we use the same training set for pre-training and fine-tuning the relation encoder. For TACRED, TACREV and ReTACRED, where the training set is comparatively smaller, we also experiment with a variant where we instead use the Gigaword corpus for pre-training the relation encoder, following the self-supervision strategy from Section 3.1. Finally, we also report results where the pre-training step is omitted and we directly train the relation encoder using the fine-tuning strategy from Section 3.2.
For our main experiments, we use roberta-large [^4] https://huggingface.co/docs/transformers/model_doc/roberta to initialise the relation encoder. We pre-train the model for 25 epochs and select the checkpoint with the minimum validation loss. For the fine-tuning step, we similarly train the model for 25 epochs and select the best checkpoint based on the validation set. We use the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of $45e-$ 4 and a weight decay $\lambda$ of $51e-$ 5. The temperature $\tau$ in the contrastive loss was set to 0.05. For the relation classifier, we used 256-dimensional embeddings for the hidden layer.
4.2 Results
TACRED
TACREV
Re-TACRED
Baselines
SpanBERT 70.8 $*{}^{*}$ ${}^{\dagger}$
LUKE 72.7 ${}^{\circ}$ ${}^{\circ}$
Typed Marker 74.6 83.2 91.1
Curriculum Learning 75.0 - 91.4
Standard pre-training
Mask 22.8 22.4 22.9
[H,T] 60.7 62.9 71.9
[H,T,Mask] 72.7 72.8 81.3
[H,T]+Mask 77.6 83.9 91.5
EnCore+Mask 77.9 83.5 91.6
EnCore+[H,T]+Mask 77.8 83.3 91.5
Gigaword pre-training
Mask 24.0 23.1 37.8
[H,T] 62.9 65.9 79.0
[H,T,Mask] 74.1 74.7 82.6
[H,T]+Mask 78.1 83.3 92.4
EnCore+Mask 78.9 84.4 93.1
EnCore+[H,T]+Mask 78.9 84.1 92.7
No pre-training
[H,T,Mask] 55.2 54.3 60.4
EnCore 69.8 76.1 80.4
[…]
Wiki-WD NYT-FB
P R F1 P@10 P@30
Baselines
RECON 87.2 87.2 87.2 87.5 74.1
KGPool 88.6 88.6 88.6 92.3 86.7
Standard pre-training
Mask 49.8 48.2 48.4 51.9 49.3
[H,T] 75.7 74.2 74.9 79.3 78.0
[H,T,Mask] 81.8 80.3 81.1 87.9 85.8
[H,T]+Mask 89.3 88.9 89.1 94.2 92.9
EnCore+Mask 89.0 88.6 88.8 94.8 92.9
EnCore+[H,T]+Mask 89.4 88.0 88.7 94.3 92.6
No pre-training
[H,T,Mask] 55.5 54.3 54.88 57.9 56.3
EnCore 80.5 78.9 79.7 85.8 84.2
Table 3: Comparison of different relation embedding strategies. Baseline results were obtained from the original papers. Our models are initialised from RoBERTa-large.
Table 2 summarises the results for TACRED, TACREV and Re-TACRED. Let us first consider the standard pre-training results (i.e. where the models are pre-trained on the standard training set). A number of observations stand out. First, the Mask strategy on its own performs poorly. Second, there is clear evidence that the [H,T] and Mask strategies are complementary: both [H,T,Mask] and [H,T]+Mask substantially outperform [H,T] and Mask on their own. We can furthermore see that [H,T]+Mask outperforms [H,T,Mask]. The combined pre-training strategy which is used by [H,T,Mask] thus indeed seems to largely fail to learn meaningful [MASK] embeddings. Finally, the EnCore+Mask strategy matches or even outperforms [H,T]+Mask. This is surprising, given that the EnCore embeddings are specifically trained to capture semantic types and are not fine-tuned on the relation extraction datasets. The strong performance of EnCore+Mask thus clearly supports the idea that the [H,T] representation mostly captures the entity types, rather than the relationship itself.
The results of the configurations that were pre-trained on Gigaword show the effectiveness of this pre-training strategy, as the results of all models are better than with standard pre-training. While the fact that pre-training on external datasets can bring benefits is in itself unsurprising, this self-supervision strategy based on coreference chains was not previously tested for relation extraction. It offers a convenient way to improve relation extraction systems, since it does not rely on an entity linked corpus. Comparing the performance of the different configurations, after pre-training on Gigaword, we see the same patterns as with standard pre-training. The Encore+Mask strategy emerges as the best model overall, which in particular confirms the usefulness of combining entity typing information with direct relation embeddings.
We can also see that forgoing pre-training altogether has a detrimental effect. A direct comparison with the baselines is difficult, as the two strongest baselines use additional information (i.e. the gold entity type labels). Nonetheless, we can see that our best configurations outperform the state-of-the-art methods. The improvements are relatively small for TACREV and Re-TACRED, especially with standard pre-training, while being somewhat larger for TACRED.
Table 3 summarises the results we obtained for Wiki-WD and NYT-FB. The main patterns are consistent with the results from Table 2. For instance, we can again see that Mask on its own performs poorly and that [H,T]+Mask outperforms both Mask and [H,T] by a considerable margin. We furthermore again see that the pre-trained entity embeddings from EnCore can serve the same purpose as the fine-tuned [H,T] embeddings. The best configurations outperform the baselines, with the improvements on NYT-FB being particularly clear. However, these methods are not directly comparable as they focus on different information.
4.3 Analysis
BERT-base
BERT-large
RoBERTa-large
ALBERT-xxl
TACRED F1 74.2 74.7 77.6 77.9
TACREV F1 74.8 76.8 82.9 83.4
Re-TACRED F1 86.0 87.4 91.5 91.9
Wiki-WD F1 79.2 84.3 89.1 89.4
NYT-FB P@10 93.7 93.9 94.2 94.5
NYT-FB P@30 93.0 92.3 92.9 93.3
Table 4: Comparison of different language models, using the [H,T]+Mask relation embedding strategy under standard pre-training.
Sentence Label
Asia Bibi was sentenced to hang in Pakistan’s central province of Punjab earlier this month after being accused of insulting the Prophet Mohammed in 2009.
Gold: no relation
[H,T]: per:state or provinces of residence
[H,T]+Mask: no relation
“To get rich is glorious" has been the mantra in booming communist China for 30 years, but few have embraced the slogan more vigorously than Wen Qiang, a leading law official in the country’s southwest.
Gold: no relation
[H,T]: per:countries of residence
[H,T]+Mask: no relation
Benjamin Chertoff is the Editor in Chief of Popular Mechanics magazine as well as the cousin of the Director of Homeland Security, Michael Chertoff.
Gold: no relation
[H,T]: per:employee of
[H,T]+Mask: no relation
WASHINGTON – The National Restaurant Association gave \$35,000 – a year’s salary – in severance pay to a female staff member in the late 1990s after an encounter with Herman Cain, its chief executive at the time, made her uncomfortable working there, three people with direct knowledge of the payment said on Tuesday.
Gold: org:top members/employees
[H,T]: no relation
[H,T]+Mask: org:top members/employees
Table 5: Comparison of the TACREV test set outputs from the [H, T] and [H, T] + Mask models that were initialised using RoBERTa-large.
Comparison of Different Language Models
In Table 4, we compare the performance of roberta-large, which we have been using for our main experiments, with three other LMs: bert-base-uncased [^5] https://huggingface.co/docs/transformers/model_doc/bert, bert-large-uncased and albert-xxlarge-v1666https://huggingface.co/docs/transformers/model_doc/albert. Unsurprisingly, we can see that bert-large-uncased outperforms bert-base-uncased. We furthermore see, across all datasets, that roberta-large outperforms bert-large-uncased, but the best results are obtained by albert-xxlarge-v1. This is consistent with the findings from Zhong and Chen (2021). The main advantage of albert-xxlarge-v1 is that it uses 4096-dimensional embeddings, compared to 1024-dimensional embeddings for bert-large-uncased and roberta-large, despite being smaller than the latter two models (due to the use of a parameter sharing across layers). This allows the model to learn more informative relation embeddings.
Qualitative Analysis of Model Outputs
In Table 5, we compare predictions of the [H,T] and [H,T]+Mask models, for examples from the TACREV test set. These examples illustrate how the [H,T] model often overly relies on the semantic types of the entities, without fully taking into account the actual sentence context. As the first two examples illustrate, given a person and a country, the [H,T] model often predicts that the person resides in that country, even if there is no relationship. The third example shows a similar issue, which arises when the sentence refers to a person and an organisation, in which case the [H,T] model often incorrectly predicts that the person is employed by that organisation. As the fourth example illustrates, however, the opposite situation also arises, where [H,T] mistakenly predicts that there is no relation. These examples support the view that the [H,T] model focuses almost exclusively on modelling the entity types, which is not always sufficient. To further illustrate this point, we present an analysis of the confusion matrices in the appendix.
5 Conclusions
The primary aim of this paper was to analyse two different strategies for training relation encoders. On the one hand, most work in supervised relation extraction relies on contextualised embeddings of the head and tail entity for predicting relationships. On the other hand, prompt-based strategies can be used to obtain embeddings that represent the relationship itself. The latter strategy is arguably more intuitive, but we found it to perform poorly in practice. Rather than suggesting that such relation embeddings are not useful, however, we found that they capture information that is highly complementary to what is captured by contextualised entity embeddings. Indeed, we considered a hybrid strategy, which substantially outperforms either of the two individual strategies, allowing us to improve the state of the art in each of the five considered benchmarks. Remarkably, we found that this remains true if we use entity embeddings from an off-the-shelf entity encoder. Finally, as a secondary contribution, we also found that coreference chains, which were used for training entity encoders in a self-supervised way by Mtumbuka and Schockaert (2023), can be successfully leveraged for self-supervised training of relation encoders.
Limitations
Our analysis in this paper was limited to the English language, and we only considered the setting of fully supervised sentence-level relation extraction. Since our focus was on learning representations (i.e. relation embeddings), it is not straightforward to transfer our findings to problems such as zero-shot or few-shot relation extraction. We did not consider the use of LLMs for relation extraction in this paper. On the one hand, this is due to the fact that applying LLMs to this setting is not straightforward, especially when the set of candidate relation labels is large. Progress in this area has indeed been slow, as we highlighted in the related work section. Moreover, while LLMs can be used to extract symbolic representations (e.g. knowledge graph triples), they are less suitable for learning embeddings. Traditionally, relation embeddings have primarily been used as an intermediate representations, before relation labels were predicted, and from this perspective, we may wonder whether relation embeddings are still needed in the LLM era. However, beyond acting as an intermediate representation, embeddings have a number of important advantages. They can, in principle, capture much more subtle distinctions than is possible with pre-defined discrete relation labels. As such, they are more suitable for modelling relational similarity (i.e. analogy), for instance.
Appendix A Additional Analysis
Comparison of Confusion Matrices
In Figures 1 and 2, we show the confusion matrix for 5 randomly sampled relations and the “no relation” label, for the TACREV test set. For this analysis, we focus on the [H,T] and [H,T]+Mask models. We group any predictions of labels outside the sampled six classes as the “other” class. We can see that the “no relation” label is the most challenging label for both models: most errors arise when the models predict a relationship when there is none, or vice versa. We can also see, however, that the [H,T]+Mask consistently improves on [H,T] when it comes to modelling the “no relation” label. This supports the conclusions from our qualitative analysis.
Figure 1: The confusion matrix for the [H, T] model on the TACREV test set.
Figure 2: The confusion matrix for the [H, T]+ Mask model on the TACREV test set.
BERT
bidirectional encoder representations from transformers
KG
knowledge graph
MLM
masked language modelling
PLM
pre-trained language model